OpenStudy (anonymous):

What is the value of n in the equation (n – 4) – 3 = 3 – (2n + 3)?

OpenStudy (anonymous):

n-4-3=3-2n-3 n-7=-2n 3n=7 n=7/3

OpenStudy (anonymous):

thanks